11. Morgan llc has base sales of 100 and depreciation expense is 10% of sales. Assuming a 10% increase in sales, depreciation expense for the first pro forma year is:

  1. 10

  2. 11

  3. 12

  4. None of the above

24. Gadson acquires digital corporation which is upstream in the marketing chain. this is an example of:

  1. Diversification

  2. Vertical integration

  3. Horizontal integration

  4. None of the above

27. Tester corporation has a beta of 2.0. the current t-bill rate is 1% and the stock market's historical return has exceeded the risk-free rate by 8%. the cost of equity for tester is:

  1. 16%

  2. 17%

  3. 18%

  4. 19%
Answer #1

Hi

As per policy we will solve only top most question here.

11) Sales = 100

depreciation = 10% of Sales

=10%*100 = 10

Sales next year = 100*(1+10%) = 110

Depreciation for pro forma year = 110*10%

= 11

hence 11 is correct option here.

Thanks